OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is odd or 1 die shows a 5? A 6 X 6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

The sum of numbers can be either even or odd.... The probability of having a sum as odd is 1/2. Probability of none of the dices showing a four is (5/6)*(5/6) = 25/36 Thus probability of at least one four is 1 - 25/36 = 11/36 Total probability = 1/2 + 11/36 = 29/36

OpenStudy (anonymous):

1/6 would be the correct answer

OpenStudy (anonymous):

nvm, it would be 23/36